OpenStudy (anonymous):

For 3x^2 - 5x - 24 identify p and q

myininaya (myininaya):

what is p and q?

OpenStudy (saifoo.khan):

the roots i guess.

OpenStudy (anonymous):

i honestly have no idea

myininaya (myininaya):

p=justin q=bieber?

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

:p q:

OpenStudy (anonymous):

3.781967 and -2.115301002

OpenStudy (dumbcow):

3*-24 = -72 ok hmm i don't think this factors

OpenStudy (anonymous):

you use the quadratic formula to solve this problem

OpenStudy (dumbcow):

weird that would ask for p and q, usually that implies it factors

myininaya (myininaya):

so jp this is all you see of the problem it just says 3x^2-5x-24 identify p and q?

OpenStudy (anonymous):

you can factor too if its possible

OpenStudy (anonymous):

yes...

myininaya (myininaya):

have you ever heard your teach repeatedly call something p and q?

myininaya (myininaya):

teacher*

OpenStudy (dumbcow):

factored form is referred to as y = (x-p)(x-q)

OpenStudy (anonymous):

no. im taking some online alg 2 course and it was one of my practice questions, and i was confused.

OpenStudy (dumbcow):

anyway, roshandapink got it use quadratic formula here

myininaya (myininaya):

perhaps p and q are the zeros it just drives me crazy that they wouldn't say identify the zeros p and q

OpenStudy (anonymous):

ok, i guess i'll look back at the section (more carefully).

OpenStudy (anonymous):

thanks anyways

OpenStudy (dumbcow):

have you used quadratic formula before

myininaya (myininaya):

\[ax^2+bx+c=0=>x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]